OpenStudy (anonymous):

log (base3) x + 13 log (base3) x^4 =14

jhonyy9 (jhonyy9):

hope you know log(a) b +log(a) c = log(a) b*c like a first step hope this will help you

jhonyy9 (jhonyy9):

and for second term 3log(a) b = log(a) b^3

jhonyy9 (jhonyy9):

so and you can rewriting the 14 like log(3) 3^14 because you know that log(3) 3 = 1

OpenStudy (anonymous):

Yeah first i moved the 13 to get log (base 3) x^17, then multipled to get log (base3) [x(x^17)]

jhonyy9 (jhonyy9):

so (x^4)^13 not is equal x^17

jhonyy9 (jhonyy9):

so (x^a)^b = x^(a*b) ok ?

OpenStudy (anonymous):

so it should be 52?

jhonyy9 (jhonyy9):

yes

OpenStudy (anonymous):

so when i combine them it would be log(base3) x^53= 14?

jhonyy9 (jhonyy9):

yes but how can you rewriting the 14 using log(3) ?

OpenStudy (anonymous):

3^14= x^54? Im not sure.

jhonyy9 (jhonyy9):

why exponent 54 ?

jhonyy9 (jhonyy9):

52+1=53

OpenStudy (anonymous):

i meant to write 53

OpenStudy (anonymous):

yeah so it would be 4782969=x^53?

jhonyy9 (jhonyy9):

you have any choice ?

OpenStudy (anonymous):

what do you mean?

jhonyy9 (jhonyy9):

what is the right result ?

OpenStudy (anonymous):

i get 115911 when i do the 53rd root of the number which it says is wrong

jhonyy9 (jhonyy9):

you dont missed anything there in this exercise ? ,because this is the right way to solve it sure

OpenStudy (anonymous):

nope im not missing anything